# Who Should I Listen To? Adaptive Collaboration in Personalized Federated Learning

### v2

@@ 19          At server at time $t$ do
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq\frac{4L\left(\mathcal{L}_{i}^{0}-\mathcal{L}_{i}^{*}\right)}{T}+\frac{\overline{\sigma}^{2}}{2Ld}+\frac{e^{2}\tilde{\sigma}^{2}}{2L|U|}+2\delta\enspace.$
The proof is provided in Appendix 6. Abourayya et al. (2025) show that under the assumption of increasing local accuracy, pseudo-labels stabilize after some round $t_{0}$, so the assumption of a bounded change in the client objective is realistic. In fact, the global loss term effectively becomes stationary under these assumptions quickly and the expected drift becomes negligibly small as $t$ increases.
Privacy considerations:
A key advantage of FedMosaic lies in its privacy-preserving design. Unlike traditional federated learning approaches that share model parameters or gradients, both known to be vulnerable to reconstruction and membership inference attacks (Zhu et al., 2019), our method shares only hard predictions on a public unlabeled dataset. This has been shown to significantly reduce privacy leakage compared to parameter- or soft-label-sharing schemes (Abourayya et al., 2025). In addition to the prediction matrix $L_{t}^{i}$, clients share a confidence vector $E_{t}^{i}$, where each element $E_{t}^{i}[j]$ quantifies their estimated expertise for the corresponding prediction. While $L_{t}^{i}$ is minimally informative, $E_{t}^{i}$ could potentially encode sensitive characteristics of the client’s private data distribution. To mitigate this risk, we propose a differentially private variant of FedMosaic (DP-FedMosaic) in which calibrated noise is added to the expertise scores before sharing. This ensures differential privacy while preserving the functional role of confidence-based collaboration. Our differentially private variant DP-FedMosaic is evaluated in sec.4. The results demonstrate that strong privacy can be guaranteed with minimal degradation in performance.
4 Empirical Evaluation
In this section, we show that FedMosaic outperforms state-of-the-art personalized federated learning methods across three challenging scenarios: (1) label skew, (2) feature shift, and (3) hybrid settings combining label and feature heterogeneity. We evaluate our method across multiple datasets and compare against strong PFL baselines, local training, and centralized upper bounds. Additional ablations and results are included in the appendix.
Label skew:
Label skew is a commonly used protocol to simulate heterogeneous output distributions, where clients observe only subsets of the label space. We consider two standard variants: (i) a pathological setting, where each client sees only a minimal subset of possible classes, and (ii) a more realistic setting where label proportions are sampled using an asymmetric distribution, inducing overlapping but imbalanced class distributions. These settings are widely adopted in the literature (T Dinh et al., 2020; Fallah et al., 2020; Zhang et al., 2023a, d, b).
In both the label skew and feature shift experiments, we use the class-frequency-based variant of the confidence score. This heuristic estimates confidence by the relative frequency of the predicted class in the local training data, which is readily available and effective when class imbalance is a primary factor of heterogeneity.
We evaluate on FashionMNIST and CIFAR10 with $m=15$ clients. In the pathological case, each client holds data from only 2 classes, resulting in extreme label imbalance. In the practical setting, label proportions are drawn from a Dirichlet distribution, simulating more natural variation. As shown in Table 2, FedMosaic consistently outperforms all baselines in both scenarios. On FashionMNIST, FedMosaic achieves $0.9940$ accuracy in the pathological setting, surpassing strong PFL methods like Ditto and APFL and even slightly outperforming both local and centralized training. On CIFAR10, FedMosaic reaches $0.8803$, again outperforming all competitors. In the practical scenario, performance trends are similar: FedMosaic achieves $0.9843$ and $0.8615$ on FashionMNIST and CIFAR10, respectively, demonstrating robustness across severity levels of label imbalance.
Table 2: Test accuracy under pathological and practical Non-IID Settings for $m=15$ clients. Color Map: baselines, worse than both baselines, worse than local training, better than both baselines.
Method Pathological non-IID Practical non-IID
FashionMNIST CIFAR10 FashionMNIST CIFAR10
Centralized $0.9928$ $(0.001)$ $0.8790$ $(0.001)$ $0.9928$ $(0.0003)$ $0.8790$ $(0.0004)$
Local training $0.9932$ $(0.0002)$ $0.8801$ $(0.0001)$ $0.9823$ $(0.0001)$ $0.8391$ $(0.002)$
FedAvg $0.7672$ $(0.001)$ $0.6442$ $(0.002)$ $0.8371$ $(0.002)$ $0.7028$ $(0.004)$
FedProx $0.7788$ $(0.003)$ $0.7025$ $(0.002)$ $0.8414$ $(0.003)$ $0.7335$ $(0.004)$
FedCT $0.7815$ $(0.0001)$ $0.7391$ $(0.0002)$ $0.8527$ $(0.0.0001)$ $0.7439$ $(0.0001)$
Per-FedAvg $0.9863$ $(0.0002)$ $0.8720$ $(0.0001)$ $0.9711$ $(0.0001)$ $0.8137$ $(0.002)$
Ditto $0.9937$ $(0.0001)$ $0.8794$ $(0.0001)$ $0.9839$ $(0.0002)$ $0.8389$ $(0.0004)$
FedBN $0.7804$ $(0.003)$ $0.8135$ $(0.005)$ $0.8539$ $(0.003)$ $0.8041$ $(0.007)$
pFedMe $0.7480$ $(0.004)$ $0.8147$ $(0.003)$ $0.8001$ $(0.001)$ $0.8161$ $(0.004)$
APFL $0.9926$ $(0.0004)$ $0.8798$ $(0.0001)$ $0.9796$ $(0.0003)$ $0.8381$ $(0.002)$
FedPHP $0.9930$ $(0.0001)$ $0.8790$ $(0.0001)$ $0.9840$ $(0.0001)$ $0.8375$ $(0.0003)$
FedMosaic $\mathbf{0.9940}$ $(0.0001)$ $\mathbf{0.8803}$ $(0.0001)$ $\mathbf{0.9843}$ $(0.0001)$ $\mathbf{0.8615}$ $(0.0001)$
Table 3: Test accuracy on Office and DomainNet datasets in feature shift scenarios: For Office-Caltech10: A, C, D ,W are abbreviations for Amazon, Caltech, DSLR and WebCam, for DomainNet: C, I, P, Q, R, S are abbreviations for Clipart, Infograph, Painting, Quickdraw, Real and Sketch. Color Map: baselines, worse than both baselines, worse than local training, better than both baselines.
Method Office DomainNet
A C D W C I P Q R S
Centralized $0.7403$ ($0.001$) $0.5824$ ($0.002$) $0.7912$ ($0.002$) $0.7852$ ($0.0001$) $0.7053$ ($0.004$) $0.3059$ ($0.003$) $0.6187$ ($0.002$) $0.7150$ ($0.001$) $0.7017$ ($0.004$) $0.6462$ ($0.003$)
Local training $0.7136$ ($0.0002$) $0.3867$ ($0.003$) $0.8125$ ($0.001$) $0.7627$ ($0.002$) $0.6531$ ($0.005$) $0.3825$ ($0.007$) $0.6652$ ($0.003$) $0.7843$ ($0.003$) $0.7104$ ($0.002$) $0.7053$ ($0.006$)
FedAvg $0.7188$ ($0.001$) $0.4844$ ($0.001$) $0.4063$ ($0.002$) $0.5424$ ($0.006$) $0.5571$ ($0.002$) $0.2842$ ($0.005$) $0.4025$ ($0.003$) $0.5264$ ($0.002$) $0.5415$ ($0.001$) $0.5612$ ($0.002$)
FedProx $0.7344$ ($0.002$) $0.5200$ ($0.002$) $0.6875$ ($0.004$) $0.7966$ ($0.004$) $0.5941$ ($0.002$) $0.3574$ ($0.004$) $0.4882$ ($0.004$) $0.5537$ ($0.001$) $0.5682$ ($0.005$) $0.5917$ ($0.002$)
FedCT $0.7396$ ($0.001$) $0.5721$ ($0.002$) $0.6873$ ($0.0001$) $0.7031$ ($0.0002$) $0.6153$ ($0.003$) $0.3519$ ($0.0001$) $0.6473$ ($0.0003$) $0.6082$ ($0.0001$) $0.7185$ ($0.0002$) $0.6925$ ($0.0001$)
Per-FedAvg $0.7304$ ($0.001$) $0.5181$ ($0.005$) $0.6922$ ($0.003$) $0.7758$ ($0.0001$) $0.6842$ ($0.0001$) $0.3621$ ($0.002$) $0.6049$ ($0.002$) $0.7263$ ($0.001$) $0.7084$ ($0.003$) $0.6816$ ($0.003$)
Ditto $0.7530$ ($0.0001$) $0.5791$ ($0.003$) $0.7839$ ($0.0002$) $0.7839$ ($0.001$) $0.7097$ ($0.0001$) $0.3913$ ($0.0001$) $0.6731$ ($0.0002$) $0.8033$ ($0.0003$) $0.7735$ ($0.0001$) $0.7314$ ($0.0003$)
FedBN $0.7539$ ($0.0001$) $0.5813$ ($0.0001$) $0.7854$ ($0.002$) $0.7823$ ($0.008$) $0.6945$ ($0.003$) $0.3801$ ($0.001$) $0.6812$ ($0.002$) $0.7921$ ($0.002$) $0.7620$ ($0.001$) $0.6923$ ($0.001$)
pFedMe $0.7083$ ($0.003$) $0.4978$ ($0.001$) $0.7500$ ($0.0003$) $0.6441$ ($0.0001$) $0.6721$ ($0.001$) $0.3742$ ($0.003$) $0.6517$ ($0.002$) $0.7524$ ($0.002$) $0.7419$ ($0.001$) $0.6893$ ($0.003$)
APFL $0.7130$ ($0.0001$) $0.3905$ ($0.0006$) $0.5085$ ($0.002$) $0.6963$ ($0.001$) $0.6873$ ($0.001$) $0.3805$ ($0.003$) $0.6739$ ($0.003$) $0.7914$ ($0.0001$) $0.7742$ ($0.001$) $0.7185$ ($0.002$)
FedPHP $0.7063$ ($0.005$) $0.4013$ ($0.0004$) $0.5178$ ($0.0001$) $0.7274$ ($0.0002$) $0.6529$ ($0.004$) $0.3632$ ($0.003$) $0.6601$ ($0.005$) $0.7703$ ($0.002$) $0.7528$ ($0.006$) $0.7011$ ($0.001$)
FedMosaic $\mathbf{0.8021}$ ($0.0001$) $\mathbf{0.6000}$ ($0.0002$) $\mathbf{0.8125}$ ($0.0002$) $\mathbf{0.8305}$ ($0.001$) $\mathbf{0.7136}$ ($0.001$) $\mathbf{0.4159}$ ($0.002$) $\mathbf{0.6938}$ ($0.004$) $\mathbf{0.8427}$ ($0.001$) $\mathbf{0.7925}$ ($0.003$) $\mathbf{0.7503}$ ($0.002$)
Feature Shift
To evaluate robustness under heterogeneous input distributions, we test FedMosaic on feature shift scenarios using Office-Caltech10 and DomainNet. In this setting, all clients share the same label space but receive data from different domains (e.g., Amazon vs. Webcam). Office-Caltech10 contains four domains (Amazon (A), Caltech(C), DSLR(D), and Webcam (W)), and DomainNet spans six domains (Clipart(C), Infograph(I), Painting(P),Quickdraw(Q),Real(R) and Sketch(S)), offering greater complexity and domain variability. While not as prominent as label skew, feature shift has scenarios have been used to evaluate FL and PFL methods Li et al. (2021c); Lee et al. (2023); Mclaughlin and Su (2024). As shown in Table 3, FedMosaic consistently achieves the highest accuracy on all domains, outperforming both traditional and personalized FL baselines. For example, on Office-Caltech10, FedMosaic attains $0.8021$ on Amazon, $0.6000$ on Caltech, $0.8125$ on DSLR, and $0.8305$ on Webcam—improving on Ditto, FedBN, and APFL. On DomainNet, FedMosaic again sets the highest scores across all six domains. These results demonstrate that our adaptive, confidence-weighted strategy effectively handles domain-specific variations while promoting generalization.
Hybrid Distribution (Label Skew + Feature Shift)
We now consider the most challenging case: combining label skew and feature shift. To simulate this hybrid heterogeneity, we partition each domain in DomainNet and Office-Caltech10 into multiple clients, assigning to each a random subset of classes. Specifically, we sample 5 clients per domain, each with access to only 2 of 10 classes (similar to the pathological label skew), resulting in 30 clients for DomainNet and 20 for Office. In this scenario, we compare both variants of our confidence-based aggregation strategy, a class-frequency-based heuristic (W) and an uncertainty-based score (U).
Table 4 shows results for multiple architectures (AlexNet and ViT). In all cases, FedMosaic significantly outperforms the best existing methods. For example, FedMosaic (U) achieves $0.8943$ accuracy on Office-Caltech10, compared to $0.8063$ for the best baseline (Ditto). On DomainNet, it reaches $0.8836$ (AlexNet) and $0.8735$ (ViT), again surpassing all alternatives. Notably, our differentially private variant (DP-FedMosaic) maintains strong performance with minimal degradation, illustrating that privacy can be ensured without compromising collaborative gains. Interestingly, both the class-frequency-based and uncertainty-based variants of FedMosaic achieve similarly strong performance across the hybrid settings. While the uncertainty-based method offers greater flexibility by capturing instance-level predictive confidence, the class-frequency heuristic performs equally well despite its simplicity. This may be due to the use of pathological label skew in our hybrid setup, where each client sees only a few classes—making frequency a strong proxy for confidence.
Taken together, these results validate that FedMosaic consistently outperforms both local and centralized baselines across diverse and realistic non-IID settings—demonstrating that adaptive, expert-aware collaboration is a principled and practical strategy for personalized federated learning.
Table 4: Test accuracy on Office and DomainNet datasets in hybrid settings scenario for $m=30$ clients on DomainNet and $m=20$ for Office dataset. Color Map: baselines, worse than both baselines, worse than local training, better than both baselines.
Method DomainNet DomainNet (ViT) Office
Centralized $0.6624$ $(0.004)$ $0.6825$ $(0.002)$ $0.4092$ $(0.006)$
Local training $0.8464$ $(0.001)$ $0.8492$ $(0.003)$ $0.8679$ $(0.004)$
FedAvg $0.3100$ $(0.008)$ $0.3328$ $(0.005)$ $0.3725$ $(0.008)$
FedProx $0.5523$ $(0.001)$ $0.5718$ $(0.003)$ $0.5839$ $(0.003)$
FedCT $0.5638$ $(0.0001)$ $0.6752$ $(0.0002)$ $0.5942$ $(0.0002)$
Per-FedAvg $0.7248$ $(0.004)$ $0.7319$ $(0.003)$ $0.7192$ $(0.005)$
Ditto $0.8147$ $(0.0001)$ $0.8382$ $(0.0002)$ $0.8063$ $(0.0001)$
FedBN $0.7154$ $(0.003)$ $0.7039$ $(0.004)$ $0.7548$ $(0.003)$
pFedMe $0.7521$ $(0.005)$ $0.7681$ $(0.008)$ $0.7483$ $(0.007)$
APFL $0.8059$ $(0.003)$ $0.8327$ $(0.005)$ $0.8091$ $(0.001)$
FedPHP $0.7825$ $(0.006)$ $0.7731$ $(0.007)$ $0.7636$ $(0.004)$
FedMosaic (W) $0.8744$ $(0.0002)$ $0.8852$ $(0.002)$ $0.8906$ $(0.0001)$
FedMosaic (U) $0.8836$ $(0.0001)$ $0.8735$ $(0.001)$ $0.8943$ $(0.0003)$
DP-FedMosaic (U) $0.8593$ $(0.001)$ $0.8574$ $(0.001)$ $0.8719$ $(0.0004)$
5 Discussion and Conclusion
Personalized Federated learning (PFL) aims to address data heterogeneity by tailoring models to client-specific distributions. Yet, as we have demonstrated, many existing approaches fall short of their promise, often failing to pass even local training or centralized baselines. This raises fundamental concerns about how collaboration is structured in personalized federated learning.
We argue that meaningful personalization in federated learning requires more than per-client modeling: it must involve adaptive, data-specific collaboration. In particular, PFL methods should support example-level decision-making, allowing clients to modulate the degree and direction of collaboration based on local context and per-sample reliability. Without this level of adaptivity, personalization risks becoming a superficial modification of global training. FedMosaic is one concrete instantiation of this principle. It enables example-level collaboration through dynamic loss weighting and confidence-based aggregation over a shared unlabeled dataset. Unlike prior methods that personalize only at the client level, FedMosaic allows each client to adapt both how much and from whom to learn, based on alignment between public and private data.
Empirical results across a diverse set of non-IID scenarios support the effectiveness of this approach. In the hybrid scenario, which combines label skew and feature shift, FedMosaic outperforms all competitors and baselines by a wide margin. In the feature shift scenarios, it again surpasses all methods across most domains, in many domains with substantial gains. In the label skew setting, FedMosaic consistently achieves the best performance for the pathological non-IID scenario, though with very narrow margins, in particular wrt. local training. In the practical non-IID scenario with milder heterogeneity, centralized training performs best, as expected. Yet, traditional federated learning methods fall short, being outperformed by several PFL approaches, including FedMosaic.
These results illustrate both strengths and limitations of personalized FL. One limitation is that, particularly in the label skew setting, the advantage over strong local baselines can be modest. Such scenarios, in particular the pathological non-IID one, raise the question of whether collaboration is justified, and whether evaluation setups with strong local baselines but weak global benefit are well-posed. We therefore emphasize the need for more meaningful benchmarks: scenarios where collaboration has clear potential upside, and where the evaluation criteria capture the practical value of federated interaction, not just statistical differences. That said, FedMosaic demonstrates that adaptive and data-aware collaboration is feasible and effective. It is the only method to outperform both local and centralized baselines across all settings, validating its robustness and practical utility.
While we provide a differentially private variant of FedMosaic, a deeper theoretical understanding of the trade-offs between privacy, collaboration, and generalization remains an open challenge. These questions go beyond our method, they point toward a more general agenda of making federated collaboration both trustworthy and adaptive. We see this work as the starting point of a broader research agenda: developing federated learning algorithms that adaptively personalize collaboration itself, not just models, and doing so in a way that is data-aware, privacy-preserving, and robust across deployment conditions.

### v3

@@ Appendix A Proof of Theorem
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq\frac{4L\left(\mathcal{L}_{i}^{0}-\mathcal{L}_{i}^{*}\right)}{T}+\frac{\overline{\sigma}^{2}}{2Ld}+\frac{e^{2}\tilde{\sigma}^{2}}{2L|U|}+2\delta\kern 5.0pt.$
Proof.
Since $\mathcal{L}_{i}^{\mathrm{local}}$ and $\mathcal{L}_{i}^{\mathrm{global},t}$ are $L(1+e)^{-1}$-smooth, and since during optimization steps $\lambda_{i}^{t}<e$ is fixed, the Lipschitz constant of $\mathcal{L}_{i}^{t}$ is
$L(1+e)^{-1}+\lambda_{i}^{t}L(1+e)^{-1}\leq L(1+e)^{-1}+eL(1+e)^{-1}=L\kern 5.0pt.$
Thus, the standard descent lemma (Bottou et al., 2018) gives:
$\mathbb{E}[\mathcal{L}_{i}^{t}(\theta_{t+1})]\leq\mathbb{E}[\mathcal{L}_{i}^{t}(\theta_{t})]-\eta\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]+\frac{L\eta^{2}}{2}\mathbb{E}[\|g_{i}^{t}\|^{2}].$
To bound $\mathbb{E}[\|g_{i}^{t}\|^{2}]$, expand
$\mathbb{E}[\|g_{i}^{t}\|^{2}]=\mathbb{E}[\|g_{i}^{t}-\nabla\mathcal{L}_{i}^{t}(\theta_{t})+\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq 2\sigma^{2}+2\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}],$
and substitute into the descent inequality to obtain
$\mathbb{E}[\mathcal{L}_{i}^{t}(\theta_{t+1})]\leq\mathbb{E}[\mathcal{L}_{i}^{t}(\theta_{t})]-\eta\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]+L\eta^{2}\left(\sigma^{2}+\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\right).$
Rearranging terms yields
$\mathbb{E}[\mathcal{L}_{i}^{t}(\theta_{t+1})]\leq\mathbb{E}[\mathcal{L}_{i}^{t}(\theta_{t})]-\eta(1-L\eta)\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]+L\eta^{2}\sigma^{2}.$
This step requires $\eta\leq(2L)^{-1}<L^{-1}$ to ensure that the coefficient $(1-L\eta)$ is positive. We now account for the fact that the function changes between rounds, i.e.,
$\mathbb{E}[\mathcal{L}_{i}^{t+1}(\theta_{t+1})]\leq\mathbb{E}[\mathcal{L}_{i}^{t}(\theta_{t+1})]+\delta,$
which gives
$\mathbb{E}[\mathcal{L}_{i}^{t+1}(\theta_{t+1})]\leq\mathbb{E}[\mathcal{L}_{i}^{t}(\theta_{t})]-\eta(1-L\eta)\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]+L\eta^{2}\sigma^{2}+\delta.$
Summing from $t=0$ to T1T-1 and rearranging yields
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq\frac{\mathcal{L}_{i}^{0}-\mathcal{L}_{i}^{T}}{(1-L\eta)\eta T}+\frac{L\eta^{2}\sigma^{2}}{1-L\eta}+\frac{\delta}{1-L\eta}.$
Denoting the minimum loss as $\mathcal{L}_{i}^{*}$, i.e., $\forall t$, $\mathcal{L}_{i}^{t}\geq\mathcal{L}_{i}^{*}$ yields the formal result
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq\frac{\mathcal{L}_{i}^{0}-\mathcal{L}_{i}^{*}}{(1-L\eta)\eta T}+\frac{L\eta^{2}\sigma^{2}}{1-L\eta}+\frac{\delta}{1-L\eta}.$
Since $((1-L\eta)\eta)^{-1}$, $L\eta^{2}/(1-L\eta)$, and $(1-L\eta)^{-1}$ have a maximum at $(2L)^{-1}$ for $\eta\leq(2L)^{-1}$, we can upper bound this by
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq\frac{4L\left(\mathcal{L}_{i}^{0}-\mathcal{L}_{i}^{*}\right)}{T}+\frac{\sigma^{2}}{4L}+2\delta\kern 5.0pt.$
Since $g_{i}^{t}=g_{i}^{local,t}+\lambda_{i}^{t}g_{i}^{global,t}$, we decompose $\sigma^{2}$ in round $t$ at client $i$ as $2\overline{\sigma}^{2}+2(\lambda_{i}^{t})^{2}\tilde{\sigma}^{2}$, and further bound
$\begin{split}\sigma_{global}^{2}\leq&\frac{\mathbb{E}_{x\in D_{i}}\left[\left\|\nabla_{\theta}\ell(\theta,x,\hat{y}^{t})-\nabla\mathcal{L}_{i}^{local,t}\right\|^{2}\right]}{\min_{i}|D_{i}|}\\ &+\sup_{i,t}(\lambda_{i}^{t})^{2}\frac{\mathbb{E}_{x\in U}\left[\left\|\nabla_{\theta}\ell(\theta,x,\hat{y}^{t})-\nabla\mathcal{L}_{i}^{global,t}\right\|^{2}\right]}{|U|}\\ \leq&\frac{2\overline{\sigma}^{2}}{d}+\frac{2e^{2}\tilde{\sigma}^{2}}{|U|}\kern 5.0pt,\end{split}$
since $\sup_{i,t}(\lambda_{i}^{t})^{2}=e^{2}$ and using Assumption 5. With this, we obtain obtain
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}[\|\nabla\mathcal{L}_{i}^{t}(\theta_{t})\|^{2}]\leq\frac{4L\left(\mathcal{L}_{i}^{0}-\mathcal{L}_{i}^{*}\right)}{T}+\frac{\overline{\sigma}^{2}}{2Ld}+\frac{e^{2}\tilde{\sigma}^{2}}{2L|U|}+2\delta\kern 5.0pt.$
Appendix B Additional Empirical Evaluation
Robustness under Misleading Global Knowledge
To further evaluate FedMosaic’s adaptivity, we conducted an experiment designed to test its behavior when the global consensus signal is actively misleading for a particular client. We constructed a scenario using CIFAR-10 dataset with $5$ clients, where client 0 was assigned flipped labels so effectively training on corrupted data. This setup results in the global pseudo labels being systematically misaligned with this client’s local distribution. As expected the client’s local model suffers a significantly higher loss when trained using the global pseudo labels compared to its own data, leading to a near zero value of $\lambda$. This confirms the intended behavior of FedMosaic: when the global signal is detrimental, the client autonomously reduces its reliance on it, effectively opting out of harmful collaboration. Fig.4 illustrates this behavior by showing the divergence between global and local loss for the corrupted client (client 0) in comparison to a non-corrupted one (client $1$). Fig. 4 shows the evolution of the adaptive weight $\lambda$ across communication rounds for all $5$ clients.
Figure 3: Local Vs Global loss across communication rounds on CIFAR-10.
Figure 4: Adaptive weight $\lambda$ across communication rounds on CIFAR-10.
Personaliztion vs. Local training In Low-collaboration Regimes
While FedMosaic consistently archives the highest accuracy across both pathological and practical label skew settings (Table2, the margin between its performance and that of local training is notably small. This observation raises a critical insight. In such scenarios, where each client’s local distribution is highly disjoint and local alignment provides limited benefit, personalization through collaboration may be unnecessary or even detrimental. Indeed, FedMosaic’s adaptive mechanism reflects this reality. The per-client weighting strategy reduces reliance on the global information when it does not align with local data. This is evident in Fig.6 and Fig.6, which show that the global loss remains consistently higher than the local loss for many clients, leading to near zero value of the adaptive weight $\lambda$ as seen in Fig.7. In such cases,FedMosaic defaults to local training behavior, effectively opting out of collaboration when it offers no advantage. This reinforces the methods’ robustness as it personalizes only when beneficial, and falls back to local training when collaboration yields little or a negative return. To ensure numerical stability in the computation of the adaptive coefficient
$\lambda_{i}^{t}=\exp\left(-\frac{\mathcal{L}_{i}^{\mathrm{global}}(\theta_{t})-\mathcal{L}_{i}^{\mathrm{local},t}(\theta_{t})}{\mathcal{L}_{i}^{\mathrm{local},t}(\theta_{t})}\right),$
, we add a small constant $\epsilon$ to the denominator to prevent division by zero.
Figure 5: Local loss across communication rounds on Fashion-MNIST for the first 66 clients.
Figure 6: Global loss across communication rounds on Fashion-MINST for the first 66 clients.
Figure 7: Adaptive weight $\lambda$ across communication rounds on Fashion-MINST for the first 66 clients.
The Effect of the Unlabeled Dataset
: As mensioned in sec.4,FedMosaic relies heavily on a shared, unlabeled public dataset $|U|$. To understand how sensitive FedMosaic is to this dataset’s characteristics, we conducted a study on CIFAR-10 dataset focusing on two critical questions: First, how does the amount of available data affect performance? Second, does it matter whether the class distribution is balanced (IID) or heavily skewed?
Impact of Public Dataset Size
: We evaluated the performance of FedMosaic using different sizes of the public unlabeled dataset, with $|U|$ set to 30003000,20002000,10001000,$500$ and $250$. For this experiment, the public dataset was always sampled in an IID fashion to ensure all classes were present. The results, summarized in Fig.8, show that the performance of FedMosaic is remarkably stable. Even as the size of the public dataset is reduced by over $90\%$ ( from 3000 to 250 samples), the drop in final test accuracy is minimal. This finding suggests that the collaboration mechanism does not require a large volume of public unlabeled data. As long as a small class-representative set of examples is available, clients can effectively share knowledge and build high-quality personalized models.
Figure 8: Test accuracy (ACC) of FedMosaic under different unlabeled dataset size $|U|$
Impact of Public Dataset Distribution
: Next, we studied the effect of the public unlabeled dataset distribution. We simulated varying degrees of distribution skew by sampling $|U|$ (with a fixed size of 3,000) using a Dirichlet distribution. We tested different values of the concentration parameter $\alpha={1,0.7,0.5,0.3,0.1}$, where $\alpha=1$ corresponds to a perfectly IID distribution and lower values induce increasingly severe skew.
As shown in Fig.10 and Fig.10, we observe a degradition in performance as the public dataset become more skewed. The most significant drop occurs at very low $\alpha$ values (e.g.,0.3,0.1 ), where some classes are absent from $U$. In such cases, the global consensus offers no useful information for clients whose private data contains the missing classes.
Figure 9: Test accuracy (ACC) of FedMosaic under different distribution of $U$.
Figure 10: Class distributiuon of $U$ under different values of alpha.
However, the most crucial finding is that the performance of FedMosaic never drops below the local training baseline. This demonstrates the robustness of the adaptive aggregation scheme. When the global signal becomes irrelevant or misleading, the dynamic loss weight $\lambda$ automatically steers clients to disregard it, effectively defaulting to local training. This acts as a critical fail-safe, ensuring that collaboration is never actively detrimental, even when the public data is of poor quality.